OpenStudy (anonymous):

Solve for n: (4^n)^3 (2^n + 2) = (4^2)^n – 1

OpenStudy (anonymous):

(n-1)!/(n-2)!+(n-2)!/(2!(n-4)!)=4 (((n-2)!)^2 + 2(n-4)! (n-1)!) / (2(n-4)! (n-2)!) = 4 1/2*(n^2 - 3n + 4) = 4 n^2 - 3n - 4 = 0 (n + 1)(n - 4) = 0 n = -1 => reject n = 4